# Caught in the act: detections of recoiling supermassive black holes from simulations

### v1

@@ 5 Discussion and Summary
From Equation 3, a more massive SMBH is more likely to host a detectable BRC than a less massive SMBH. Consequently, the prime targets for investigation are those galaxies which are expected to host the most massive SMBHs, namely cored galaxies (e.g. Thomas et al., 2016). A large stellar core in an ETG is most readily attributed to SMBH binary scouring (e.g. Thomas et al., 2014): observing a core indicates that a SMBH binary was likely present in the galaxy at some point, and hence a recoil candidate is plausible. Assuming a stellar core radius of the order $1\,\mathrm{kpc}$, corresponding to an angular size of $\sim 0\farcs 06$ at $z=0.6$, massive cored ETGs below this redshift are an ideal first place to look for BRCs with current instrument capability.
A key result of our study is the dependence of BRC observability on recoil velocity (Figure 3). We found a recoil velocity of $v_{\mathrm{kick}}\sim 540\,\mathrm{km}\,\mathrm{s}^{-1}$ to represent a particularly challenging velocity to detect: mass of the BRC is less than that for lower recoil velocities, and the low SMBH $r_{\mathrm{apo}}$ reduces the contrast of the BRC compared to the host galaxy. We expect that for different initial conditions, the most challenging BRC to observe would correspond to $r_{\mathrm{d}}(v_{\mathrm{kick}})\simeq R_{\mathrm{e}}$, however this would most likely be in the vicinity of $v_{\mathrm{kick}}\sim 600\,\mathrm{km}\,\mathrm{s}^{-1}$ given most massive ETGs have $R_{\mathrm{e}}\sim 6.0\,\mathrm{kpc}$ (e.g. Kelvin et al., 2012), similar to our models with $R_{\mathrm{e}}\sim 5.6\,\mathrm{kpc}$. Comparatively, the most likely BRCs to be detected are those which have $v_{\mathrm{kick}}>\sigma_{\star,0}$ and $r_{\mathrm{d}}(v_{\mathrm{kick}})<R_{\mathrm{e}}$: these kicks are more frequent than $v_{\mathrm{kick}}\simeq 0.5v_{\mathrm{esc}}$, and are accompanied by a BRC massive enough for detection.
@@
This work has focused on detections of recoiling SMBHs by way of their enshrouding stellar mass, however the techniques presented here may also pave an avenue for detecting wandering black holes that have been ejected from lower mass systems, either dynamically or due to GW-induced recoil (e.g. Rizzuto et al., 2022; Rantala et al., 2024a; Rantala & Naab, 2025). Wandering black holes formed in nuclear stellar clusters are also expected to posess an entourage of bound stellar material, however as these black holes are generally of lower mass than those considered in this work, one would expect their stellar velocity dispersions to be lower, of the order a few tens to a hundred $\mathrm{km}\,\mathrm{s}^{-1}$ (e.g. Stone et al., 2017).
@@
where $\frac{\mathrm{d}N_{\mathrm{m}}}{\mathrm{d}\xi\mathrm{d}z}(M,\xi,z)$ is fit to the collisionless Millennium and Millennium II simulations in Fakhouri et al. (2010). Multiplying this result with the cumulative halo mass function for $M_{\mathrm{DM}}\gtrsim 5\times 10^{13}\,\mathrm{M}_{\sun}$ (e.g. Rodríguez-Puebla et al., 2016) results in a $\sim 2\times 10^{-4}\,\mathrm{Mpc}^{-3}$ major merger density. Hence, out to a redshift $z=1$ where the comoving volume is $\sim 1.7\times 10^{11}\,\mathrm{Mpc}^{3}$, there are $\sim 4\times 10^{7}$ mergers satisfying our constraints, and a corresponding number of recoiling SMBHs. Repeating the exercise for $z=0.6$ results in a value of $\sim 7\times 10^{6}$ recoiling SMBHs. We estimated $\sim 20\%$ of recoiling SMBHs would at some point in their excursion from the galaxy centre be observable (subsection 3.2); this translates to $\sim 8\times 10^{6}$ BRCs out to $z=1$. Estimates for how long such BRCs would be observable is significantly more complicated, as torquing of the SMBH away from the initial recoiling axis can increase the amount of time the SMBH is below the detection limit $r_{\mathrm{d}}(v_{\mathrm{kick}})$ in projection. Clearly, the largest window of observability occurs for purely radial motion of the settling SMBH, and is thus most probable up until first apocentre, which we found to have a median value of $t_{\mathrm{apo}}\sim 30\,\mathrm{Myr}$. Thus, when considering the observable duration of a BRC, we may roughly estimate the number of detectable BRCs to be:
@@
To provide robust estimates for the total duration for which a BRC would be observable, thus refining the estimate given by Equation 9, one would need to extend the study performed here to include a variety of galaxy triaxialities, obtained through cosmological simulations where the impact of the surrounding environment could be quantified. Additionally, one might consider the effect stochastic properties of the SMBH binary prior to coalescence might have on the BRC, and hence its detectability. For example, a low SMBH binary eccentricity arising from small phase space perturbations (e.g. Rawlings et al., 2023) would increase the overall time spent in the binary scouring phase before GW-driven coalescence, lowering the central density (Rantala et al., 2024b) and potentially also the amount of stellar mass available to form the BRC.
@@ 6 Acknowledgments
We thank Till Sawala and Maximilian Fabricius for useful discussions. A.R. acknowledges the support by the University of Helsinki Research Foundation. A.R. and P.H.J. acknowledge the support by the European Research Council via ERC Consolidator Grant KETJU (no. 818930). P.H.J also acknowledges the support of the Research Council of Finland grant 339127. T.N. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311 from the DFG Cluster of Excellence “ORIGINS”. This research has made use of the SVO Filter Profile Service “Carlos Rodrigo”, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00. We list here the roles and contributions of the authors according to the Contributor Roles Taxonomy (CRediT). AR: conceptualisation, methodology, formal analysis, investigation, data curation, writing: original. PHJ: supervision, writing: original, funding acquisition. TN: conceptualisation, investigation, writing: original. AR: writing: review. JT: writing: review. BN: writing: review.
@@ Appendix B Line of sight velocity distribution
To generate the mock IFU images, pseudo-particles are divided using a Voronoi tesselation. The LOS velocity distribution of pseudo-particles within each Voronoi bin are then fit using a Gauss-Hermite series (after fitting for the mean velocity $V$ and velocity dispersion $\sigma$) defined (van der Marel & Franx, 1993):

### v2

@@ 6 Discussion and Summary
From Equation 3, a more massive SMBH is more likely to host a detectable BRC than a less massive SMBH. Consequently, the prime targets for investigation are those galaxies which are expected to host the most massive SMBHs, namely cored galaxies (e.g. thomas2016). A large stellar core in an ETG is most readily attributed to SMBH binary scouring (e.g. Thomas2014): observing a core indicates that a SMBH binary was likely present in the galaxy at some point, and hence a recoil candidate is plausible. Assuming a stellar core radius of the order $1\,\mathrm{kpc}$, corresponding to an angular size of $\sim 0\farcs 06$ at $z=0.6$, massive cored ETGs below this redshift are an ideal first place to look for BRCs with current instrument capability.
A key result of our study is the dependence of BRC observability on recoil velocity (Figure 3). We found a recoil velocity of $v_{\mathrm{kick}}\sim 540\,\mathrm{km}\,\mathrm{s}^{-1}$ to represent a particularly challenging velocity to detect: mass of the BRC is less than that for lower recoil velocities, and the low SMBH $r_{\mathrm{apo}}$ reduces the contrast of the BRC compared to the host galaxy. We expect that for different initial conditions, the most challenging BRC to observe would correspond to $r_{\mathrm{d}}(v_{\mathrm{kick}})\simeq R_{\mathrm{e}}$, however this would most likely be in the vicinity of $v_{\mathrm{kick}}\sim 600\,\mathrm{km}\,\mathrm{s}^{-1}$ given most massive ETGs have $R_{\mathrm{e}}\sim 6.0\,\mathrm{kpc}$ (e.g. kelvin2012), similar to our models with $R_{\mathrm{e}}\sim 5.6\,\mathrm{kpc}$. Comparatively, the most likely BRCs to be detected are those which have $v_{\mathrm{kick}}>\sigma_{\star,0}$ and $r_{\mathrm{d}}(v_{\mathrm{kick}})<R_{\mathrm{e}}$: these kicks are more frequent than $v_{\mathrm{kick}}\simeq 0.5v_{\mathrm{esc}}$, and are accompanied by a BRC massive enough for detection.
@@
This work has focused on detections of recoiling SMBHs by way of their enshrouding stellar mass, however the techniques presented here may also pave an avenue for detecting wandering black holes that have been ejected from lower mass systems, either dynamically or due to GW-induced recoil (e.g. rizzuto2022; rantala2024b; rantala2025). Wandering black holes formed in nuclear stellar clusters are also expected to posess an entourage of bound stellar material, however as these black holes are generally of lower mass than those considered in this work, one would expect their stellar velocity dispersions to be lower, of the order a few tens to a hundred $\mathrm{km}\,\mathrm{s}^{-1}$ (e.g. stone2017).
@@
where $\frac{\mathrm{d}N_{\mathrm{m}}}{\mathrm{d}\xi\mathrm{d}z}(M,\xi,z)$ is fit to the collisionless Millennium and Millennium II simulations in fakhouri2010. Multiplying this result with the cumulative halo mass function for $M_{\mathrm{DM}}\gtrsim 5\times 10^{13}\,\mathrm{M}_{\sun}$ (e.g. rodriguezpuebla2016) results in a $\sim 2\times 10^{-4}\,\mathrm{Mpc}^{-3}$ major merger density. Hence, out to a redshift $z=1$ where the comoving volume is $\sim 1.7\times 10^{11}\,\mathrm{Mpc}^{3}$, there are $\sim 4\times 10^{7}$ mergers satisfying our constraints, and a corresponding number of recoiling SMBHs. Repeating the exercise for $z=0.6$ results in a value of $\sim 7\times 10^{6}$ recoiling SMBHs. We estimated $\sim 20\%$ of recoiling SMBHs would at some point in their excursion from the galaxy centre be observable (subsection 3.2); this translates to $\sim 8\times 10^{6}$ BRCs out to $z=1$. Estimates for how long such BRCs would be observable is significantly more complicated, as torquing of the SMBH away from the initial recoiling axis can increase the amount of time the SMBH is below the detection limit $r_{\mathrm{d}}(v_{\mathrm{kick}})$ in projection. Clearly, the largest window of observability occurs for purely radial motion of the settling SMBH, and is thus most probable up until first apocentre, which we found to have a median value of $t_{\mathrm{apo}}\sim 30\,\mathrm{Myr}$. Thus, when considering the observable duration of a BRC, we may roughly estimate the number of detectable BRCs to be:
@@
To provide robust estimates for the total duration for which a BRC would be observable, thus refining the estimate given by Equation 9, one would need to extend the study performed here to include a variety of galaxy triaxialities (our merger remnant has a triaxiality of $T=\left(a^{2}-b^{2}\right)/\left(a^{2}-c^{2}\right)\simeq 0.82$), obtained through cosmological simulations where the impact of the surrounding environment could be quantified. Additionally, one might consider the effect stochastic properties of the SMBH binary prior to coalescence might have on the BRC, and hence its detectability. For example, a low SMBH binary eccentricity arising from small phase space perturbations (e.g. rawlings2023) would increase the overall time spent in the binary scouring phase before GW-driven coalescence, lowering the central density (rantala2024) and potentially also the amount of stellar mass available to form the BRC.
@@ 7 Acknowledgments
We thank the anonymous referee for their comments which helped improve the manuscript. We also thank Till Sawala and Maximilian Fabricius for useful discussions. A.R. acknowledges the support by the University of Helsinki Research Foundation. A.R. and P.H.J. acknowledge the support by the European Research Council via ERC Consolidator Grant KETJU (no. 818930). P.H.J also acknowledges the support of the Research Council of Finland grant 339127. T.N. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311 from the DFG Cluster of Excellence “ORIGINS”. This research has made use of the SVO Filter Profile Service “Carlos Rodrigo”, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00. We list here the roles and contributions of the authors according to the Contributor Roles Taxonomy (CRediT). AR: conceptualisation, methodology, formal analysis, investigation, data curation, writing: original. PHJ: supervision, writing: original, funding acquisition. TN: conceptualisation, investigation, writing: original. AR: writing: review. JT: writing: review. BN: writing: review.
@@ Appendix B Line of sight velocity distribution
To generate the mock IFU images, pseudo-particles are divided using a Voronoi tesselation. The LOS velocity distribution of pseudo-particles within each Voronoi bin are then fit using a Gauss-Hermite series (after fitting for the mean velocity $V$ and velocity dispersion $\sigma$) defined (vandermarel1993):
@@
Appendix C IFU comparison between major and minor mergers
Figure 8: Comparison between major mergers (top two rows) and minor mergers (bottom two rows) in velocity dispersion $\sigma$ (left column) and first order asymmetric deviation $h_{3}$ (right column). The times are normalised such that $t=0$ corresponds to the time of SMBH binary coalescence. The major merger has a recoiling SMBH with $v_{\mathrm{kick}}=540\,\mathrm{km}\,\mathrm{s}^{-1}$, and is the same IFU maps as in the middle centre and middle right panels of Figure 4. Clear asymmetric lobes in velocity dispersion are present, but no strong spatial correlation between Voronoi pixels in $h_{3}$ are visible. The minor merger has no recoiling SMBH, with the seen velocity dispersion asymmetry at $t=0.05,\mathrm{Gyr}$ arising from the asymmetry in the merger. At this moment there is strong spatial correlation in $h_{3}$ at small radii, with evidence for two kinematically-distinct regions. As the merger remnant settles, there is a symmetric velocity dispersion, reminiscent of the no-recoil equal-mass merger in Figure 4. The $h_{3}$ map has reduced spatial correlation.
We show the mock IFU maps for the $v_{\mathrm{kick}}=540\,\mathrm{km}\,\mathrm{s}^{-1}$, equal-mass merger remnant at two moments following SMBH coalescence, and similar maps for the no recoil unequal-mass merger, in Figure 8. We see that whilst unequal-mass mergers can produce asymmetric velocity dispersion features, as do recoiling SMBHS, the former also produces coherent features in the first order asymmetric deviation coefficient $h_{3}$, which a recoiling SMBH does not.